# Vector Field-Guided Learning Predictive Control for Motion Planning of Mobile Robots with Uncertain Dynamics

### v2

@@ 38 Function Main():
The optimal solution $\tilde{u}^{*}(\hat{\tilde{x}}_{\tau})$ and the optimal value function $\lambda^{*}(\hat{\tilde{x}}_{\tau})$ are difficult to be solved analytically. This is because $\lambda^{*}(\hat{\tilde{x}}_{\tau+1})$ is not available; see Eqs. (31) and (32). However, they can be approximated by two kernel-based networks, i.e.,
$\hat{\tilde{u}}(\hat{\tilde{x}}_{\tau})={W}_{a,\tau}^{\top}\Phi(\hat{\tilde{x}}_{\tau}),$ (33a)
$\hat{\lambda}(\hat{\tilde{x}}_{\tau})={W}_{c,\tau}^{\top}\Phi(\hat{\tilde{x}}_{\tau}),$ (33b)
@@
where $c_{1},\cdots,c_{n_{\Phi}}$ are the elements in the RL training dictionary and $k(\cdot,\cdot):\mathbb{R}^{n_{\mathcal{K}}}\times\mathbb{R}^{n_{\mathcal{K}}}\rightarrow\mathbb{R}$ is the Gaussian kernel function.
With the estimated value function in (33b) and consistent with (32), we can obtain the target optimal solution at the $\tau$-th horizon by
$\tilde{u}(\hat{\tilde{x}}_{\tau})=-\dfrac{1}{2}\gamma R^{-1}B_{d,\tau}^{\top}\hat{\lambda}(\hat{\tilde{x}}_{\tau+1}).$ (34)
@@
$\lambda(\hat{\tilde{x}}_{\tau})=\left\{\begin{array}[]{cc}2Q\hat{\tilde{x}}_{\tau}+\gamma A_{d,\tau}^{\top}\hat{\lambda}(\hat{\tilde{x}}_{\tau+1})+\dfrac{\partial h(\hat{x}_{\tau})}{\partial\hat{\tilde{x}}_{\tau}},&\tau\in[k,k+N-1]\\ 2P\hat{\tilde{x}}_{k+N}+{\partial h({\hat{x}}_{k+N})}/{\partial\hat{\tilde{x}}_{k+N}},&\tau=k+N\end{array}\right.$
Next, we present the iterative learning processes. From the previous formulations, $N$ actor-critic networks are to be constructed to approximate the optimal value function and solution in (31) and (32). At the $i$-th iteration, the critic network aims at minimizing the error function between the target and the approximate value functions, i.e.,
@@
${W}_{c,\tau+1}^{i}$ $={W}_{c,\tau}^{i}-\eta_{c}{\partial\epsilon_{c}^{i}(\hat{\tilde{x}}_{\tau})}/{\partial W_{c,\tau}^{i}},$ (38a)
${W}_{a,\tau+1}^{i}$ $={W}_{a,\tau}^{i}-\eta_{a}{\partial\epsilon_{a}^{i}(\hat{\tilde{x}}_{\tau})}/{\partial W_{a,\tau}^{i}},$ (38b)
where $\eta_{c},\eta_{a}>0$ are the step coefficients of the gradients. At each time instant, the initial weights $A_{\tau}$ and $C_{\tau}$ inherit from the $(\tau+1)$-th weights of the previous time instant, and $A_{N}$ and $C_{N}$ inherit from the $(N-1)$-th weights.
The overall algorithm of VF-LPC is shown in Algorithm. 1. Also, the convergence theorem of the Eb-RHRL algorithm is given below and its proof is presented in Section V.
@@
(Convergence of the Eb-RHRL). As the iteration number $i$ increases, sequences $\lambda^{i}$ and $\tilde{u}^{i}$ converge to (30) and (32) respectively, i.e., as $i\rightarrow\infty$, $\lambda^{i}\rightarrow\lambda^{*}$ and $\tilde{u}^{i}\rightarrow\tilde{u}^{*}$.
@@ V Theoretic Analysis
This section presents the theoretical analyses regarding Theorem 1. We prove that the Eb-RHRL algorithm, integrating the game-based barrier function, converges to the optimal value function and solution, thereby solving the OMP problem. Proving the convergence of the Eb-RHRL algorithm is equivalent to prove $\lambda^{i}\rightarrow\lambda^{*}$ and $\tilde{u}^{i}\rightarrow\tilde{u}^{*}$ as $i\rightarrow\infty$.
@@
Given a control sequence $\{\mu^{i}\}$ involving random actions and taking (24) into (29), one can obtain the following equation:
$\Lambda^{i+1}(\hat{\tilde{x}}_{\tau})=\hat{\tilde{x}}_{\tau}^{\top}Q\hat{\tilde{x}}_{\tau}+\mu^{i}(\hat{\tilde{x}}_{\tau})^{\top}R\mu^{i}(\hat{\tilde{x}}_{\tau})+h(\hat{x}_{\tau})+\Lambda^{i}(\hat{\tilde{x}}_{\tau+1}),$
where $\mu(\cdot):\mathbb{R}^{n_{\mathcal{K}}}\rightarrow\mathbb{R}^{n_{u}}$. Rewrite (29) as follows:
$V^{i+1}(\hat{\tilde{x}}_{\tau})=\hat{\tilde{x}}_{\tau}^{\top}Q\hat{\tilde{x}}_{\tau}+\tilde{u}^{i}(\hat{\tilde{x}}_{\tau})^{\top}R\tilde{u}^{i}(\hat{\tilde{x}}_{\tau})+h(\hat{x}_{\tau})+V^{i}(\hat{\tilde{x}}_{\tau+1}),$
@@
Lemma 1 shows that applying $\{\tilde{u}^{i}\}$ to the robot system results in a lower value function than applying a random solution sequence. It is now used to obtain and prove Lemma 2.
@@
Let the sequence $V^{i}(\hat{\tilde{x}}_{\tau})$ be defined as in (28). By applying an initial admissible control policy, the following conclusions hold: 1) There exists an upper bound $\bar{Z}(\hat{\tilde{x}}_{\tau})$ such that $0\leq V^{i}(\hat{\tilde{x}}_{\tau})\leq\bar{Z}(\hat{\tilde{x}}_{\tau})$ for all $i$. 2) If (27) is solvable, then there exists a upper bound $\bar{Z}(\hat{\tilde{x}}_{\tau})$ such that $V(\hat{\tilde{x}}_{\tau})\leq\bar{Z}(\hat{\tilde{x}}_{\tau})$, where $V(\hat{\tilde{x}}_{\tau})$ solves (26) and $0\leq V^{i}(\hat{\tilde{x}}_{\tau})\leq V^{*}(\hat{\tilde{x}}_{\tau})\leq\bar{Z}(\hat{\tilde{x}}_{\tau})$ for all $i$.
@@
To begin with, $\eta(\cdot):\mathbb{R}^{n_{\mathcal{K}}}\rightarrow\mathbb{R}^{n_{u}}$ is assumed to be an admissible policy. Let $V^{0}(\hat{\tilde{x}}_{\tau})=Z^{0}(\hat{\tilde{x}}_{\tau})=0$, where
$Z^{i+1}(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\eta(\hat{\tilde{x}}_{\tau}))+Z^{i}(\hat{\tilde{x}}_{\tau+1})\right]$
with $Z^{k+N}(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left(F(\hat{\tilde{x}}_{k+N}\right)$, for all $i\geq 1$. We define an upper bound
$\bar{Z}(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left[\sum_{j=\tau}^{k+N-1}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))+F(\hat{\tilde{x}}_{k+N})\right],$
with the dynamic model: $\hat{\tilde{x}}_{\tau+1}=A_{d,\tau}\hat{\tilde{x}}_{\tau}+B_{d,\tau}\eta(\hat{\tilde{x}}_{\tau})$, where $L(\hat{\tilde{x}}_{\tau})=\hat{\tilde{x}}_{\tau}^{\top}Q\hat{\tilde{x}}_{\tau}+\eta^{\top}(\hat{\tilde{x}}_{\tau})R\eta(\hat{\tilde{x}}_{\tau})+h(\hat{x}_{\tau})$. With (28), the following two cases are analyzed, respectively.
Case 1 ($\tau+i\leq k+N-1$):
$Z^{i+1}(\hat{\tilde{x}}_{\tau})$ $=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\eta(\hat{\tilde{x}}_{\tau}))+Z^{i}(\hat{\tilde{x}}_{\tau+1})\right]$ (39)
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+1}L\left(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j})\right)+Z^{i-1}(\hat{\tilde{x}}_{\tau+2})\right]$
@@
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+i}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))\right]$
Since $h(\hat{x}_{j})\in\{0,\mu_{o}\}$, we have $L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))\geq 0$. For $i_{1},i_{2}\in[k,k+N-1]$, one has
$\sum_{j=\tau}^{\tau+i_{1}}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))\leq\sum_{j=\tau}^{\tau+i_{2}}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j})),i_{1}\leq i_{2}.$ (40)
Combining Eqs. (39) and (40), we have
$Z^{i+1}(\hat{\tilde{x}}_{\tau})$ (41)
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+i}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))\right]$
$\leq\mathbb{E}\left[\sum_{j=\tau}^{k+N-1}L(\hat{\tilde{x}}_{j},\eta_{j}(\hat{\tilde{x}}_{j}))+F(\hat{\tilde{x}}_{k+N})\right]=\bar{Z}(\hat{\tilde{x}}_{\tau}).$
Case 2 ($\tau+i>k+N-1$):
@@
Combining Eqs. (41) and (42), $Z^{i+1}(\hat{\tilde{x}}_{\tau})\leq\bar{Z}(\hat{\tilde{x}}_{\tau})$ holds.
[…]

### v4

@@ 38 Function Main():
At the $i$-th iteration, the two optimal policies of the actor and critic networks are approximated by two kernel-based-network structures, i.e.,
$\hat{\tilde{u}}^{i}(\hat{\tilde{x}}_{\tau})$ $=({W}_{a,\tau}^{i})^{\top}\Phi(\hat{\tilde{x}}_{\tau}),$ (33a)
$\hat{\lambda}^{i}(\hat{\tilde{x}}_{\tau})$ $=({W}_{c,\tau}^{i})^{\top}\Phi(\hat{\tilde{x}}_{\tau}),$ (33b)
@@
where $c_{1},\cdots,c_{n_{\Phi}}$ are elements in the RL training dictionary and $k(\cdot,\cdot):\mathbb{R}^{n_{\mathcal{K}}}\times\mathbb{R}^{n_{\mathcal{K}}}\rightarrow\mathbb{R}$ is the Gaussian kernel function.
With the estimated value function in (33b) and consistent with (32), we can obtain the target optimal solution at the $i$-th iteration and the $\tau$-th horizon by
$\tilde{u}^{i}(\hat{\tilde{x}}_{\tau})=-\dfrac{1}{2}\gamma R^{-1}B_{d,\tau}^{\top}\hat{\lambda}^{i}(\hat{\tilde{x}}_{\tau+1}),\tau\in[k,k+N-1].$ (34)
Correspondingly, for $\tau\in[k,k+N]$, the value function $V^{i+1}(\hat{\tilde{x}}_{\tau})$ in Eq. (28) is written by
$V^{i+1}(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\tilde{u}^{i}(\hat{\tilde{x}}_{\tau}))+\gamma V^{i}(\hat{\tilde{x}}_{\tau+1})\right].$ (35)
@@
$\lambda^{i+1}(\hat{\tilde{x}}_{\tau})=\left\{\begin{array}[]{cc}2Q\hat{\tilde{x}}_{\tau}+\gamma A_{d,\tau}^{\top}\hat{\lambda}^{i}(\hat{\tilde{x}}_{\tau+1})+\dfrac{\partial h(\hat{x}_{\tau})}{\partial\hat{\tilde{x}}_{\tau}},&\tau\in[k,k+N-1]\\ 2P\hat{\tilde{x}}_{k+N}+{\partial h({\hat{x}}_{k+N})}/{\partial\hat{\tilde{x}}_{k+N}},&\tau=k+N\end{array}\right$.
The critic network aims at minimizing the error function between the target and the approximate value functions, i.e.,
@@
${W}_{c,\tau}^{i+1}$ $={W}_{c,\tau}^{i}-\eta_{c}{\partial\epsilon_{c}^{i}(\hat{\tilde{x}}_{\tau})}/{\partial W_{c,\tau}^{i}},$ (39a)
${W}_{a,\tau}^{i+1}$ $={W}_{a,\tau}^{i}-\eta_{a}{\partial\epsilon_{a}^{i}(\hat{\tilde{x}}_{\tau})}/{\partial W_{a,\tau}^{i}},$ (39b)
where $\eta_{c},\eta_{a}>0$ are the step coefficients of the gradients.
The overall algorithm of VF-LPC is shown in Algorithm. 1. Also, the convergence theorem of the Eb-RHRL algorithm is given below and its proof is presented in Section V. In the following, expressions similar to $\{\lambda^{i}\}$ and $\{\tilde{u}^{i}\}$ denote the sequences in the prediction horizon of the $k$-th time instant; i.e.,
$\{\lambda^{i}\}:=(\lambda^{i}(\hat{\tilde{x}}_{k}),\cdots,\lambda^{i}(\hat{\tilde{x}}_{k+N-1})),$
and
$\{\tilde{u}^{i}\}:=(\tilde{u}^{i}(\hat{\tilde{x}}_{k}),\cdots,\tilde{u}^{i}(\hat{\tilde{x}}_{k+N-1})).$
@@
(Convergence of the Eb-RHRL). As the iteration number $i$ increases, sequences $\{\lambda^{i}\}$ and $\{\tilde{u}^{i}\}$ converge to (31) and (32), respectively; i.e., $\lim_{i\rightarrow\infty}\{\lambda^{i}\}=\{\lambda^{*}\}$ and $\lim_{i\rightarrow\infty}\{\tilde{u}^{i}\}=\{\tilde{u}^{*}\}$.
@@ V Theoretic Analysis
This section presents the theoretical analyses regarding Theorem 1. Building upon the proofs in [38, 39, 11], we prove that the Eb-RHRL algorithm, integrating the game-based barrier function, converges to the optimal value function and solution, thereby solving the OMP problem.
@@
Given a control sequence $\{\mu^{i}\}$ involving random actions, $\Lambda^{i+1}(\hat{\tilde{x}}_{\tau})$ is defined by
$\Lambda^{i+1}(\hat{\tilde{x}}_{\tau})=\hat{\tilde{x}}_{\tau}^{\top}Q\hat{\tilde{x}}_{\tau}+\mu^{i}(\hat{\tilde{x}}_{\tau})^{\top}R\mu^{i}(\hat{\tilde{x}}_{\tau})+h(\hat{x}_{\tau})+\gamma\Lambda^{i}(\hat{\tilde{x}}_{\tau+1}),$
where $\mu^{i}(\cdot):\mathbb{R}^{n_{\mathcal{K}}}\rightarrow\mathbb{R}^{n_{u}}$. Rewrite (35) as follows:
$V^{i+1}(\hat{\tilde{x}}_{\tau})=\hat{\tilde{x}}_{\tau}^{\top}Q\hat{\tilde{x}}_{\tau}+\tilde{u}^{i}(\hat{\tilde{x}}_{\tau})^{\top}R\tilde{u}^{i}(\hat{\tilde{x}}_{\tau})+h(\hat{x}_{\tau})+\gamma V^{i}(\hat{\tilde{x}}_{\tau+1}),$
@@
Lemma 1 is now used to obtain and prove Lemma 2.
@@
Let the sequence $\{V^{i}\}$ be defined by (35). By applying an initial admissible control policy, the following conclusions hold: 1) There exists an upper bound $\bar{Z}^{i}(\hat{\tilde{x}}_{\tau})$ such that $0\leq V^{i}(\hat{\tilde{x}}_{\tau})\leq\bar{Z}^{i}(\hat{\tilde{x}}_{\tau})$ for all $i$. 2) If (32) is solvable, then there exists an upper bound $\bar{Z}^{i}(\hat{\tilde{x}}_{\tau})$ such that $V^{i}(\hat{\tilde{x}}_{\tau})\leq\bar{Z}^{i}(\hat{\tilde{x}}_{\tau})$, where $V^{i}(\hat{\tilde{x}}_{\tau})$ solves (30) and $0\leq V^{i}(\hat{\tilde{x}}_{\tau})\leq V^{*}(\hat{\tilde{x}}_{\tau})\leq\bar{Z}^{i}(\hat{\tilde{x}}_{\tau})$ for all $i$.
@@
To begin with, $\eta^{i}(\cdot):\mathbb{R}^{n_{\mathcal{K}}}\rightarrow\mathbb{R}^{n_{u}}$ is assumed to be an admissible policy. Let $V^{0}(\hat{\tilde{x}}_{\tau})=Z^{0}(\hat{\tilde{x}}_{\tau})=0$, where
$Z^{i+1}(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\eta^{i}(\hat{\tilde{x}}_{\tau}))+\gamma Z^{i}(\hat{\tilde{x}}_{\tau+1})\right],\forall i\geq 1.$
We define an upper bound
$\bar{Z}^{i}(\hat{\tilde{x}}_{\tau})$ $=\mathbb{E}\left[\sum_{j=\tau}^{\tau+N-1}\gamma^{j-\tau}L(\hat{\tilde{x}}_{j},\eta^{i}(\hat{\tilde{x}}_{j}))\right$. (40)
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \left.+\gamma^{N}Z^{i-(N-1)}(\hat{\tilde{x}}_{\tau+N})\right]$
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+N-1}\gamma^{j-\tau}L(\hat{\tilde{x}}_{j},\eta^{i}(\hat{\tilde{x}}_{j}))+\gamma^{N}F(\hat{\tilde{x}}_{\tau+N})\right]$
with the dynamic model (22) and $L(\hat{\tilde{x}}_{j},\eta^{i}(\hat{\tilde{x}}_{j}))=\hat{\tilde{x}}_{j}^{\top}Q\hat{\tilde{x}}_{j}+(\eta^{i}(\hat{\tilde{x}}_{j}))^{\top}R\eta^{i}(\hat{\tilde{x}}_{j})+h(\hat{x}_{j})$. When propagating $Z^{i+1}(\hat{\tilde{x}}_{\tau})$ from current stage $\tau$ to $\tau+N$, two cases exist and analyzed below.
Case 1 ($i\leq N-1$):
$Z^{i+1}(\hat{\tilde{x}}_{\tau})$ $=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\eta^{i}(\hat{\tilde{x}}_{\tau}))+\gamma Z^{i}(\hat{\tilde{x}}_{\tau+1})\right]$ (41)
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+1}\gamma^{j-\tau}L\left(\hat{\tilde{x}}_{j},\eta^{i}(\hat{\tilde{x}}_{j})\right)+\gamma^{2}Z^{i-1}(\hat{\tilde{x}}_{\tau+2})\right]$
@@
$=\mathbb{E}\left[\sum_{j=\tau}^{\tau+i}\gamma^{j-\tau}L(\hat{\tilde{x}}_{j},\eta^{i}(\hat{\tilde{x}}_{j}))\right].$
Since $h(\hat{x}_{j})\geq 0$, by applying an admissible control $\eta^{i}(\cdot)$ to the system, we can obtain
$\sum_{j=\tau}^{\tau+i_{1}}\gamma^{j-\tau}L(\hat{\tilde{x}}_{j},\eta^{i}(\hat{\tilde{x}}_{j}))<\sum_{j=\tau}^{\tau+i_{2}}\gamma^{j-\tau}L(\hat{\tilde{x}}_{j},\eta^{i}(\hat{\tilde{x}}_{j})),$ (42)
where $i_{1}<i_{2}\in[0,N]$. With Eqs. (41) and (42), we have
$\begin{aligned} &Z^{i+1}(\hat{\tilde{x}}_{\tau})\\ &=\mathbb{E}\left[\sum_{j=\tau}^{\tau+i}\gamma^{j-\tau}L(\hat{\tilde{x}}_{j},\eta^{i}(\hat{\tilde{x}}_{j}))\right]\\ &<\mathbb{E}\left[\sum_{j=\tau}^{\tau+N-1}\gamma^{j-\tau}L(\hat{\tilde{x}}_{j},\eta^{i}(\hat{\tilde{x}}_{j}))+\gamma^{N}F(\hat{\tilde{x}}_{\tau+N})\right]\\ &=\bar{Z}^{i}(\hat{\tilde{x}}_{\tau}).\end{aligned}$
Case 2 ($i\geq N$):
$\begin{aligned} &Z^{i+1}(\hat{\tilde{x}}_{\tau})\\ &=\mathbb{E}\left[\sum_{j=\tau}^{\tau+N-1}\gamma^{j-\tau}L(\hat{\tilde{x}}_{j},\eta^{i}(\hat{\tilde{x}}_{j}))+\gamma^{N}F(\hat{\tilde{x}}_{\tau+N})\right]=\bar{Z}^{i}(\hat{\tilde{x}}_{\tau}).\end{aligned}$
Combining Eqs. (43) and (44), $Z^{i+1}(\hat{\tilde{x}}_{\tau})\leq\bar{Z}^{i}(\hat{\tilde{x}}_{\tau})$ holds.
@@
Lemma 3.
[11, Monotonicity property] Let $V^{i}(\hat{\tilde{x}}_{\tau})$ be defined by (35). With $V^{0}(\hat{\tilde{x}}_{\tau})=0$, $V^{i+1}(\hat{\tilde{x}}_{\tau})\geq V^{i}(\hat{\tilde{x}}_{\tau})$ holds.
Proof.
The proof is given in[38].
@@
Theorem 2.
(Stability of the VF-LPC). The closed-loop system (19) is stable.
Proof.
Define the Lyapunov candidate function $W(\tau):=V(\hat{\tilde{x}}_{\tau})$ for the closed-loop system (19). One has
$W(\tau)=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\tilde{u}(\hat{\tilde{x}}_{\tau}))+V(\hat{\tilde{x}}_{\tau+1})\right].$
It can be deduced that the Lyapunov function equals zero at the origin of the system state and remains strictly positive at all non-origin points. In the following, we will prove that $\Delta W=W(\tau+1)-W(\tau)<0,\forall\tau\in[k,k+N-1]$. For $k\leq\tau<k+N-1$, $W(\tau+1)-W(\tau)\leq-\mathbb{E}[L(\hat{\tilde{x}}_{\tau},\tilde{u}(\hat{\tilde{x}}_{\tau}))]$. For $\tau=k+N-1$, one can set a function $F(\cdot)$ such that the inequality $\mathbb{E}\left[W(\tau)\right]=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\tilde{u}(\hat{\tilde{x}}_{\tau}))+V(\hat{\tilde{x}}_{\tau+1})\right]\leq\mathbb{E}\left[F(\hat{\tilde{x}}_{\tau})\right]$ holds, thereby deriving $W(\tau+1)=\mathbb{E}(L(\hat{\tilde{x}}_{\tau+1},\tilde{u}(\hat{\tilde{x}}_{\tau+1})))\leq F(\hat{\tilde{x}}_{\tau+1})=W(\tau)-L(\hat{\tilde{x}}_{\tau},\tilde{u}(\hat{\tilde{x}}_{\tau}))$. Since $L(\hat{\tilde{x}}_{\tau},\tilde{u}(\hat{\tilde{x}}_{\tau}))\geq 0$ and $L(\hat{\tilde{x}}_{\tau},\tilde{u}(\hat{\tilde{x}}_{\tau}))\geq 0$, it is shown that Lyapunov function exhibits a decreasing behavior in the prediction horizon.
In the following, we will demonstrate that the value function of the system (19) at two distinct time instants decreases monotonically. Let $\bar{L}(\hat{\tilde{x}}_{k+N})=L(\hat{\tilde{x}}_{k+N},K\hat{\tilde{x}}_{k+N})$, where $K$ is defined as the feedback gain matrix. For the brevity of analysis, define $V(\tilde{x}_{k},N)=\sum_{\tau=k}^{k+N-1}L(\hat{\tilde{x}}_{\tau},u(\tau))+V_{f}(\hat{\tilde{x}}_{k+N})$. Now one could set proper terminal value functions to make the condition $V(\hat{\tilde{x}}_{k+N+1})-V(\hat{\tilde{x}}_{k+N})\leq-L(\hat{\tilde{x}}_{k+N})$ hold. Then we have
$V^{*}(\hat{\tilde{x}}_{k+1},N)$ (45)
$=\sum_{\tau=k+1}^{k+N}L\left(\hat{\tilde{x}}_{\tau},u^{*}(\hat{\tilde{x}}_{\tau})\right)+V_{f}(\hat{\tilde{x}}_{k+N+1})$
$\leq V(\hat{\tilde{x}}_{k+1},N)$
$=\sum_{\tau=k+1}^{k+N-1}L\left(\hat{\tilde{x}}_{\tau},\tilde{u}^{*}(\hat{\tilde{x}}_{\tau})\right)+\bar{L}(\hat{\tilde{x}}_{k+N})+V_{f}(\hat{\tilde{x}}_{k+N+1})$
$=V^{*}(\tilde{x}_{k},N)-L\left(\tilde{x}_{k},\tilde{u}^{*}(\tilde{x}_{k})\right)+\bar{L}(\hat{\tilde{x}}_{k+N})$
$\,\,\,\,\,\,\,+\left(V_{f}(\hat{\tilde{x}}_{k+N+1})-V_{f}(\hat{\tilde{x}}_{k+N})\right).$
As $\bar{L}(\hat{\tilde{x}}_{k+N})\geq 0$, one has
$V_{f}(\hat{\tilde{x}}_{k+N+1})-V_{f}(\hat{\tilde{x}}_{k+N})$ (46)
$\leq-\bar{L}(\hat{\tilde{x}}_{k+N})$
$=-L^{\top}(\hat{\tilde{x}}_{k+N})\left(Q+K^{\top}RK\right)L(\hat{\tilde{x}}_{k+N}).$
Substituting Eq. (46) into Eq. (45) yields
$V^{*}(\hat{\tilde{x}}_{k+1},N)\leq V^{*}(\tilde{x}_{k},N)-L(\tilde{x}_{k},\tilde{u}^{*}(\tilde{x}_{k})).$ (47)
Therefore, we can conclude that $V^{*}(\tilde{x}_{k},N)$ is monotonically decreasing over time $k$. One has
$V^{*}(\hat{\tilde{x}}_{\infty},N)\leq V^{*}(\tilde{x}_{k},N)-\sum_{\tau=k}^{\infty}L\left(\hat{\tilde{x}}_{\tau},\tilde{u}^{*}(\hat{\tilde{x}}_{\tau})\right)$ (48)
and $\forall k\geq 0,V^{*}(\tilde{x}_{k},N)\geq 0$. It can be obtained that when $k\rightarrow\infty$, we have $L\left(\tilde{x}_{k},\tilde{u}^{*}(\tilde{x}_{k})\right)\rightarrow 0$. Moreover, since $Q,R$ are positive definite matrices, it follows that $\tilde{x}_{k}\rightarrow\mathbf{0},u^{*}(\tilde{x}_{k})\rightarrow\mathbf{0}$, as $k\rightarrow\infty$. Therefore, system (19) is asymptotically stable. When the modeling error is sufficiently small, the original system (2) is also asymptotically stable.